OpenStudy (anonymous):

A 6 meter ladder is leaning on a vertical wall. The foot of the ladder is being pulled at a rate of 0.15m/s. Calculate the rate of which the height of the ladder is decreasing when the ladder is 3 meters off the ground. I started by doing dHdT=dHdR∗dRdT

OpenStudy (anonymous):

It didn't copy correctly. \[\frac{ dH }{ dT }= \frac{ dH }{ dR }*\frac{ dR }{ dT }\]

OpenStudy (anonymous):

So first draw a picture of it ^_^

OpenStudy (anonymous):

You can start this problem a few ways, but I think the most intuitive is with labeling the ladder/wall/floor system as a right triangle.

OpenStudy (anonymous):

Also, before we get started, what're you defining as R?

OpenStudy (anonymous):

Okay Im back sorry.

OpenStudy (anonymous):



OpenStudy (anonymous):

Cool. Now in your equation you were using H - which side is it?

OpenStudy (anonymous):

H is the wall. R is my 0.15m/s

OpenStudy (anonymous):

To use the rate like that doesn't actually do anything in this case. If we label the picture with one more variable, B, - one that represents the base of the triangle  - then the rate is actually \[rate = \frac{dB}{dt} = .15m/s\]